# Probabilistic Modeling of Antibody Kinetics Post Infection and Vaccination: A Markov Chain Approach

### v1

@@ Abstract
Understanding the dynamics of antibody levels is crucial for characterizing the time-dependent response to immune events: either infections or vaccinations. The sequence and timing of these events significantly influence antibody level changes. Despite extensive interest in the topic in the recent years and many experimental studies, the effect of immune event sequences on antibody levels is not well understood. Moreover, disease or vaccination prevalence in the population are time-dependent. This, alongside the complexities of personal antibody kinetics, makes it difficult to analyze a sample immune measurement from a population. As a solution, we design a rigorous mathematical characterization in terms of a time-inhomogeneous Markov chain model for event-to-event transitions coupled with a probabilistic framework for the post-event antibody kinetics of multiple immune events. We demonstrate that this is an ideal model for immune event sequences, referred to as personal trajectories. This novel modeling framework surpasses the susceptible-infected-recovered (SIR) characterizations by rigorously tracking the probability distribution of population antibody response across time. To illustrate our ideas, we apply our mathematical framework to longitudinal severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) data from individuals with multiple documented infection and vaccination events. Our work is an important step towards a comprehensive understanding of antibody kinetics that could lead to an effective way to analyze the protective power of natural immunity or vaccination, predict missed immune events at an individual level, and inform booster timing recommendations.
@@ 1 Introduction
While understanding individual antibody kinetics is essential, it is equally important to analyze the population-level antibody distribution to inform public health decisions. Random sampling of the population for antibody testing and analyzing the information from such testing can yield valuable guidance for making population level decisions [5, 6]. The population level antibody distribution is governed by the prevalence for disease versus vaccination. Many seroprevalence studies have been conducted (e.g., [7, 8, 9]), and a few mathematical approaches to seroprevalence analysis were developed, including, for example, estimation of the force of infection [10]. The complex interplay of different factors affecting population antibody distributions suggests great value in developing a coherent mathematical framework to analyze measurements in support of decision-making.
@@
No prior mathematical theory has simultaneously addressed (i) antibody levels, (ii) prevalence, (iii) multiple classes, (iv) time-dependence, and (v) multiple immune events, nor their complex multi-scale interactions, to the best of our knowledge. In one example topics (iii) and (iv) were analyzed, but the effects of each event was studied separately. In this work, the authors applied models based on exponential and power-law kinetics to the post-peak immune vaccination response, studying several types of antibodies and investigating distributions for halving times and model parameters [11]. Standard models for disease transmission, such as susceptible-infected-recovered (SIR) models or statistical regression models, can be employed to study topics (ii)–(v) but cannot track population-level antibody responses across time (e.g., [12, 13, 14, 15]). In [16], the authors designed an age-structured model for boosting and waning immunity from infection or vaccination to severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) infection, thus addressing topics (ii)-(v). In another transmission model example, a network model was designed to characterize the spread of multiple contagions through a population (ii), (iii) [17]. A recent within-host differential equation model investigated reinfection (v) and considered the increased immune capacity induced by an additional infection with a variant [18]; such models ignore population-level trends. One research group [19] used Markov chain Monte Carlo methods to study multiple influenza infections or vaccinations in ferrets, addressing antibody levels (i) and time-dependence (iv). Limited theory has attempted to characterize the antibody response over time to infection and/or vaccination beyond our prior work [20, 21]; in the latter we studied the setting in which individuals can be infected or vaccinated but not both, and designed an unbiased prevalence estimation scheme via transition probability matrices.
@@
Our aim in this paper is to address prevalence and time-dependence in a multiclass and multi-event setting. We begin by constructing the general framework of the problem in Section 2. Next, we introduce personal response models for antibody response after an arbitrary number of immune events in Section 3. We then present a two-event time-inhomogeneous model (Section 4), expanding the number of allowed classes from the five in our previous work [20] to thirteen. The corresponding analysis is assembled in parallel to our previous work. The aforementioned model provides insight into the most general time-inhomogeneous model, which is introduced in Section 5. The corresponding conditional probabilities can be written in terms of the time-dependent antibody responses and the probability of a particular immune event sequence. This is the most mathematically complete model, but corresponding estimations can be intractable in the presence of low amounts of data. However, provided the incidence rates change slowly or have stabilized, we can utilize a time-homogeneous model to approximate the transmission. Section 5.3 considers a time-homogeneous model with an unlimited number of allowable immune events, with incidence rates assumed to be constant over time. In Section 6, we demonstrate the results of Section 4 numerically by building antibody response models for multiple events using data from two separate cohorts reporting multiple immune events, first reported in [22] and [23], and [24] [^1] Certain commercial equipment, instruments, software, or material are identified in this paper in order to specify the experimental procedure adequately. Such identification is not intended to imply recommendation or endorsement by the National Institute of Standards and Technology, nor is it intended to imply that the materials or equipment identified are necessarily the best available for the purpose.. Section 7 demonstrates the results of Section 5.3 numerically by simulating population transmission. The discussion includes further analysis of the reduced time-inhomogeneous model, limitations, extensions, and implications for immunologists (Section 8).
@@ 2 Modeling paradigm: broad ideas
We study antibody kinetics in a population as a disease spreads, with implications for public health policy decisions. Our goal is thus to assign meaning to a set of antibody measurements randomly sampled from a population on a given day. Some associated terms and definitions from applied diagnostics are listed below. In the next subsection, we describe notation specific to this paper.
@@
  • The vaccinated class consists of individuals whose most recent immune event is immunization against a disease.
@@ 2.1 Modeling framework
The states we consider are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected ($\mathcal{I}^{\prime}$), and previously vaccinated ($\mathcal{V}^{\prime}$). In the next section, we expand from these five states to thirteen to count secondary immune events as explicit, separate events. Each state has an associated prevalence: $q_{N}(T),q_{\mathcal{I}}(T),q_{\mathcal{V}}(T),q_{\mathcal{I}^{\prime}}(T)$, and $q_{\mathcal{V}^{\prime}}(T)$, denoting the fraction of the population in each state at time step $T$. Prevalence quantifies the total fraction of the population incident into that state so far and thus takes values in the range $[0,1]$.
A related quantity is the incidence rate, or fraction of the total population that newly moves to a particular state at time step $T$ of the absolute timeline; these quantities also take values in the range $[0,1]$. The incidences describe the movement from a particular state to another; $f_{\mathcal{I}N}$ describes the incidence from the naïve state to the newly infected state. Among the incidence rates, notice that $f_{\mathcal{I}N}(T),$ for example, means that this is the fraction of the population as a whole who moved from naïve in the previous step $T-1$ to newly infected now at time $T$. In other words, $f_{\mathcal{I}N}+f_{\mathcal{I}\mathcal{I}^{\prime}}+f_{\mathcal{I}\mathcal{V}^{\prime}}=f_{\mathcal{I}}$. This $f_{\mathcal{I}}(T)$ represents the fraction of the population that has moved into the newly infected class on time step $T$.
A transition matrix $S$ contains the probabilities of moving between states. Here, $S(i,j)$ is the probability of moving to state $i$ from state $j$. Let $T=0$ index the emergence of a disease. We assume that our initial state vector is $X_{1}=\bm{e}_{1}$ to model the disease emergence, so that all the members of the population are in state $N$ with probability $1$ on the day before the disease emerges. Here, $\bm{e}_{1}$ is the first unit vector. This assumption simplifies our model: we characterize an emergent disease rather than an endemic virus. Let $X_{j}$ denote the state, or class, at time step $j$. Denote the transition probabilities by $s$.
@@ 3 Personal response models
Each slice of the model $R_{z}(r,t)$ in personal timeline $t$ is a probability distribution that evolves in time.
@@
We are now able to extend our model to deal with multiple events, significantly increasing its usefulness, as this is frequently what one encounters in real-life scenarios. Let $\{z_{m}\}$ be the sequence of $\mathcal{M}$ events for an individual, with $\bm{\tau}$ being the vector of corresponding gap between subsequent events. Then, the modified shape function is
@@
Notice that for every additional immune event, we have to only estimate three additional parameters, $\{\theta_{z_{\mathcal{M}}}\}$, $\{\phi_{z_{\mathcal{M}}}\}$, $\{k_{z_{\mathcal{M}}}\}$, improving the usability of the model. As before, the $\theta$ and $\phi$ values for the first $\mathcal{M}-1$ events is already determined by the MLE for the previous models. We will require that $\{k_{z_{\mathcal{M}}}\}$ take on only one of two values, one for infection, $k_{\mathcal{I}}$, and the other for vaccination, $k_{\mathcal{V}}$, if data dictates that the decay rate is always the same following a particular immune event. We emphasize that this general response model is the convolution of the responses for the sequence of events for an individual. Now, we assess two different models to study the transmission of infection and vaccination through the population and how their responses convolve with the said transmission to generate the probability density of the population response on an absolute timeline.
@@ 4 Two-event time-inhomogeneous model
In this section, we present a model in which two immune events are allowable: two infections, two vaccinations, or one after the other. This is a generalization of our work in [21] in the sense that to incorporate multiple events, we consider new states instead of letting individuals pass between the original five states. The states that we consider for the two-event transmission model described in this section are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected once ($\mathcal{I}^{\prime}$), previously vaccinated once ($\mathcal{V}^{\prime}$), newly infected with a second infection ($\mathcal{I}\mathcal{I}^{\prime}$), newly vaccinated after infection ($\mathcal{V}\mathcal{I}^{\prime}$), newly infected after vaccination ($\mathcal{I}\mathcal{V}^{\prime}$), newly vaccinated with a second vaccination ($\mathcal{V}\mathcal{V}^{\prime}$), previously infected twice ($\mathcal{I}^{\prime}\mathcal{I}^{\prime}$), previously infected and vaccinated ($\mathcal{I}^{\prime}\mathcal{V}^{\prime}$), previously vaccinated and infected ($\mathcal{V}^{\prime}\mathcal{I}^{\prime}$), and previously vaccinated twice ($\mathcal{V}^{\prime}\mathcal{V}^{\prime}$). We employ a graph to represent our framework, in which each state or class is a node and transitions between classes are directed, weighted edges. We let $s_{N}(T)$ denote the weight of the degenerate edge to $N$ from $N$, or the probability of staying naïve. The probabilities $s_{\mathcal{I}N}(T)$ and $s_{\mathcal{V}N}(T)$ weight the edges to $\mathcal{I}$ from $N$ and to $\mathcal{V}$ from $N$, indicating infection or vaccination, respectively; the other transition probabilities have analogous interpretations. Fig. 1 shows the allowable movements between the thirteen states. Note, once an individual has experienced two immune events, they move into either a previously infected or previously vaccinated state with probability $1$ and then stay there indefinitely. In this model, we also forbid explicit movement back to the naïve state after an immune event has occurred; doing so assumes that the time scale of the problem is short.
@@ 4.1 Transition probabilities
The transition matrix for movement from time step $T-1$ to time step $T$ is given by (11). Here, the ordering of the matrix rows and columns is $N,\mathcal{I},\mathcal{V},\mathcal{I}^{\prime},\mathcal{V}^{\prime},\mathcal{I}\mathcal{I}^{\prime},\mathcal{V}\mathcal{I}^{\prime},\mathcal{I}\mathcal{V}^{\prime},\mathcal{V}\mathcal{V}^{\prime},\mathcal{I}^{\prime}\mathcal{I}^{\prime},\mathcal{V}^{\prime}\mathcal{I}^{\prime},\mathcal{I}^{\prime}\mathcal{V}^{\prime},\mathcal{V}^{\prime}\mathcal{V}^{\prime}$. In Fig. 1, this corresponds to an ordering of top to bottom and left to right within a level.
@@
Here, $\sigma-t$ is used to enforce indexing of the product in the correct order. Assuming everyone starts in the naïve class on time step 0, $H_{\sigma}\mathbf{e}_{1}$ is a $(13\times 1)$ vector with the $j^{th}$ entry corresponding to the probability that an arbitrary individual in the population ends up in the $j^{th}$ state in $\sigma$ time steps. The matrix multiplication to obtain $H_{\sigma}$ accounts for every sequence that ends in a particular immune event. See Appendix A for explicit calculations, and specifically (A3). As an example, consider the following,
@@
$\text{Prob}(X_{T}=\mathcal{I}^{\prime})$ is partitioned by the day of infection, which can be any day $t$ from 0 to $T-1$; this movement is indicated by the transition probability $s_{\mathcal{I}N}(t)$. Here, one is naïve before the day of infection, moves with probability 1 to the previously infected state $\mathcal{I}^{\prime}$ on the next day (day $t+1$), and then stays previously infected thereafter up to and including day $T$. Note that $\text{Prob}(X_{T}=N)=q_{N}(T)$; that is, the naïve state probability is equal to the corresponding prevalence. Similarly, $\text{Prob}(X_{T}=\mathcal{I})=f_{\mathcal{I}}(T)$ and $\text{Prob}(X_{T}=\mathcal{V})=f_{\mathcal{V}}(T)$, so the single new infection and single new vaccination state probabilities are equal to the corresponding incidences.

### v2

@@ Abstract
Understanding dynamics of antibody levels is crucial for characterizing time-dependent response to immune events: either infections or vaccinations. The sequence and timing of these events significantly influence antibody level changes. Despite extensive interest in the recent years and many experimental studies, the effect of immune event sequences on antibody levels is not well understood. Moreover, disease or vaccination prevalence in the population are time-dependent. This, alongside the complexities of personal antibody kinetics, makes it arduous to analyze a sample immune measurement from a population. A rigorous mathematical characterization can inform public health decision making.
Relevance to Life Sciences. A key result of this paper is an antibody response modeling framework for an arbitrary number of multiclass immune events–the first of its kind to the best of our knowledge. Our model is ideal for characterizing immune event sequences, referred to as personal trajectories. To illustrate our ideas, we apply our mathematical framework to longitudinal severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) data from individuals with multiple documented infection and vaccination events. This approach is fully generalizable to other diseases that exhibit waning immunity, such as influenza, respiratory syncytial virus (RSV), and pertussis. Our work is an important step towards a comprehensive understanding of antibody kinetics for infectious diseases that could lead to an effective way to analyze the protective power of natural immunity or vaccination, predict missed immune events at an individual level, and inform booster timing recommendations.
Mathematical Content. We design a rigorous mathematical characterization in terms of a time-inhomogeneous Markov chain model for event-to-event transitions coupled with a probabilistic framework for the post-event antibody kinetics of multiple immune events. Probabilistic models appropriately describe these measurements as they capture the natural variability in a population’s antibody response. We build probability density models for population response since the emergence of a disease via a discrete convolution of immune state transmission probabilities and personal response models, repeatedly invoking the definition of conditional probability and the law of total probability. Importantly, our coupled framework simultaneously tracks immune state and antibody response. This novel modeling approach surpasses the susceptible-infected-recovered (SIR) characterizations by rigorously tracing the probability distribution of population antibody response across time.
@@ 1 Introduction
While understanding individual antibody kinetics is essential, it is equally important to analyze the population-level antibody distribution to inform public health decisions. Random population sampling for antibody testing and analyzing such test information can yield valuable guidance for making population-level decisions [5, 6]. The population-level antibody distribution is governed by the prevalence for disease versus vaccination. Many seroprevalence studies have been conducted (e.g., [7, 8, 9]), and a few mathematical approaches to seroprevalence analysis were developed, including estimation of the force of infection [10]. The complex interplay of different factors affecting population antibody distributions suggests great value in developing a coherent mathematical framework to analyze measurements in support of decision-making. No prior mathematical theory has simultaneously addressed (i) antibody levels, (ii) prevalence, (iii) multiple classes, (iv) time-dependence, and (v) multiple immune events, nor their complex multi-scale interactions, to the best of our knowledge. In one example topics (iii) and (iv) were analyzed, but the effects of each event was studied separately. In this work, the authors applied models based on exponential and power-law kinetics to the post-peak immune vaccination response, studying several types of antibodies and investigating distributions for halving times and model parameters [11]. Standard models for disease transmission, such as susceptible-infected-recovered (SIR) models or statistical regression models, can be employed to study topics (ii)–(v) but cannot track population-level antibody responses across time (e.g., [12, 13, 14, 15]). In [16], the authors designed an age-structured model for boosting and waning immunity from infection or vaccination to severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) infection, thus addressing topics (ii)-(v). In another transmission model example, a network model was designed to characterize the spread of multiple contagions through a population (ii), (iii) [17]. A recent within-host differential equation model investigated reinfection (v) and considered the increased immune capacity induced by an additional infection with a variant [18]; such models ignore population-level trends. One research group [19] used Markov chain Monte Carlo methods to study multiple influenza infections or vaccinations in ferrets, addressing antibody levels (i) and time-dependence (iv). Limited theory has attempted to characterize the antibody response over time to infection and/or vaccination beyond our prior work [20, 21]; in the latter we studied the setting in which individuals can be infected or vaccinated but not both, and designed an unbiased prevalence estimation scheme via transition probability matrices.
@@
Our aim in this paper is to address prevalence and time-dependence in a multiclass and multi-event setting. We begin by constructing the general framework of the problem in Section 2. Next, we introduce personal response models for antibody response after an arbitrary number of immune events in Section 3. We then present a two-event time-inhomogeneous model (Section 4), expanding the number of allowed classes from the five in our previous work [20] to thirteen. The corresponding analysis is assembled in parallel to our previous work. The aforementioned model provides insight into the most general time-inhomogeneous model, which is introduced in Section 5. The corresponding conditional probabilities can be written in terms of the time-dependent antibody responses and the probability of a particular immune event sequence. This is the most mathematically complete model, but corresponding estimations can be intractable in the presence of low amounts of data. However, provided the incidence rates change slowly or have stabilized, we can utilize a time-homogeneous model to approximate the transmission. In Section 5.3, we consider a time-homogeneous model with an unlimited number of allowable immune events, with incidence rates assumed to be constant over time. Next, in Section 6 we demonstrate the results of Section 4 numerically by building antibody response models for multiple events using data from two separate cohorts reporting multiple immune events, first reported in [22], [23], and [24] [^1] Certain commercial equipment, instruments, software, or material are identified in this paper in order to specify the experimental procedure adequately. Such identification is not intended to imply recommendation or endorsement by the National Institute of Standards and Technology, nor is it intended to imply that the materials or equipment identified are necessarily the best available for the purpose.. Our Section 7 demonstrates the results of Section 5.3 numerically by simulating population transmission. The discussion includes further analysis of the reduced time-inhomogeneous model, limitations, extensions, and implications for immunologists (Section 8). Section 9 includes relevant mathematical background for non-mathematicians.
@@ 2 Modeling paradigm: broad ideas
We study antibody kinetics in a population as a disease spreads, with implications for public health policy decisions. Our goal is to assign meaning to a set of antibody measurements randomly sampled from a population on a given day. Below, we list relevant terms and definitions from applied diagnostics. In the next subsection, we describe notation specific to this paper.
@@
  • The vaccinated class consists of individuals whose most recent immune event is immunization.
@@ 2.1 Modeling framework
The states we consider are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected ($\mathcal{I}^{\prime}$), and previously vaccinated ($\mathcal{V}^{\prime}$). In the next section, we expand from these five states to thirteen to count secondary immune events as explicit, separate events. Each state has an associated prevalence: $q_{N}(T),q_{\mathcal{I}}(T),q_{\mathcal{V}}(T),q_{\mathcal{I}^{\prime}}(T)$, and $q_{\mathcal{V}^{\prime}}(T)$, denoting the fraction of the population in each state at time step $T$. Prevalence quantifies the total fraction of the population incident into that state so far and thus takes values in the range [0,1][0,1][ 0 , 1 ].
A related quantity is the incidence rate, or fraction of the total population that newly moves to a particular state at time step $T$ of the absolute timeline; these quantities also take values in the range [0,1][0,1][ 0 , 1 ]. The incidences describe the movement from a particular state to another; $f_{\mathcal{I}N}$ describes the incidence from the naïve state to the newly infected state. Among the incidence rates, notice that $f_{\mathcal{I}N}(T),$ for example, means that this is the fraction of the population as a whole who moved from naïve in the previous step $T-1$ to newly infected now at time $T$. In other words, $f_{\mathcal{I}N}+f_{\mathcal{I}\mathcal{I}^{\prime}}+f_{\mathcal{I}\mathcal{V}^{\prime}}=f_{\mathcal{I}}$. This $f_{\mathcal{I}}(T)$ represents the fraction of the population that has moved into the newly infected class on time step $T$.
A transition matrix $S$ contains the probabilities of moving between states. Here, $S(i,j)$ is the probability of moving to state $i$ from state $j$. Let $T=0$ index the emergence of a disease. We assume that our initial state vector is $X_{1}=\bm{e}_{1}$ to model the disease emergence, so that all the members of the population are in state $N$ with probability 111 on the day before the disease emerges. Here, $\bm{e}_{1}$ is the first unit vector. This assumption simplifies our model: we characterize an emergent disease rather than an endemic virus. Let $X_{j}$ denote the state, or class, at time step $j$. Denote the transition probabilities by $s$.
@@ 3 Personal response models
Each slice of the model $R_{z}(r,t)$ in personal timeline $t$ is a probability distribution; $R_{z}(r,t)$ evolves in time.
@@
We now extend our model to consider multiple events, significantly increasing its usefulness; this situation is frequently encountered in real-life scenarios. Let $\{z_{m}\}$ be the sequence of $\mathcal{M}$ events for an individual, with $\bm{\tau}$ the vector of corresponding time gaps between subsequent events. Then, the modified shape function is
@@
Notice that for every additional immune event, we must only estimate three additional parameters, $\{\theta_{z_{\mathcal{M}}}\}$, $\{\phi_{z_{\mathcal{M}}}\}$, $\{k_{z_{\mathcal{M}}}\}$, improving the usability of the model. As before, the $\theta$ and $\phi$ values for the first $\mathcal{M}-1$ events are already determined by the MLE for the previous models. If data dictates a consistent decay rate following a particular immune event, we will require that $\{k_{z_{\mathcal{M}}}\}$ take on only one of two values, one for infection, $k_{\mathcal{I}}$, and the other for vaccination, $k_{\mathcal{V}}$. We emphasize that this general response model is the convolution of the responses for the sequence of events for an individual. Now, we assess two different models to study the transmission of infection and vaccination through the population and how their responses convolve with said transmission to generate the probability density of the population response on an absolute timeline.
@@ 4 Two-event time-inhomogeneous model
In this section, we present a model in which two immune events are allowable: two infections, two vaccinations, or one after the other. This is a generalization of our work in [21] in the sense that to incorporate multiple events, we consider new states instead of letting individuals pass between the original five states. The states that we consider for the two-event transmission model described in this section are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected once ($\mathcal{I}^{\prime}$), previously vaccinated once ($\mathcal{V}^{\prime}$), newly infected with a second infection ($\mathcal{I}\mathcal{I}^{\prime}$), newly vaccinated after infection ($\mathcal{V}\mathcal{I}^{\prime}$), newly infected after vaccination ($\mathcal{I}\mathcal{V}^{\prime}$), newly vaccinated with a second vaccination ($\mathcal{V}\mathcal{V}^{\prime}$), previously infected twice ($\mathcal{I}^{\prime}\mathcal{I}^{\prime}$), previously infected and vaccinated ($\mathcal{I}^{\prime}\mathcal{V}^{\prime}$), previously vaccinated and infected ($\mathcal{V}^{\prime}\mathcal{I}^{\prime}$), and previously vaccinated twice ($\mathcal{V}^{\prime}\mathcal{V}^{\prime}$). We employ a graph to represent our framework, in which each state or class is a node and transitions between classes are directed, weighted edges. We let $s_{N}(T)$ denote the weight of the degenerate edge to $N$ from $N$, or the probability of staying naïve. The probabilities $s_{\mathcal{I}N}(T)$ and $s_{\mathcal{V}N}(T)$ weight the edges to $\mathcal{I}$ from $N$ and to $\mathcal{V}$ from $N$, indicating infection or vaccination, respectively; the other transition probabilities have analogous interpretations. Figure 1 shows the allowable movements between the thirteen states. Note, once an individual has experienced two immune events, they move into either a previously infected or previously vaccinated state with probability 111 and then stay there indefinitely. In this model, we also forbid explicit movement back to the naïve state after an immune event has occurred; doing so assumes that the time scale of the problem is short.
@@ 4.1 Transition probabilities
The transition matrix for movement from time step $T-1$ to time step $T$ is given by (11). Here, the ordering of the matrix rows and columns is $N,\mathcal{I},\mathcal{V},\mathcal{I}^{\prime},\mathcal{V}^{\prime},\mathcal{I}\mathcal{I}^{\prime},\mathcal{V}\mathcal{I}^{\prime},\mathcal{I}\mathcal{V}^{\prime},\mathcal{V}\mathcal{V}^{\prime},\mathcal{I}^{\prime}\mathcal{I}^{\prime},\mathcal{V}^{\prime}\mathcal{I}^{\prime},\mathcal{I}^{\prime}\mathcal{V}^{\prime},\mathcal{V}^{\prime}\mathcal{V}^{\prime}$. In Figure 1, this corresponds to an ordering of top to bottom and left to right within a level.
@@
Here, $\sigma-t$ is used to enforce indexing of the product in the correct order. Assuming everyone starts in the naïve class on time step 0, $H_{\sigma}\mathbf{e}_{1}$ is a (13×1)(13$\times$ 1)( 13 × 1 ) vector with the $j^{th}$ entry corresponding to the probability that an arbitrary individual in the population ends up in the $j^{th}$ state in $\sigma$ time steps. The matrix multiplication to obtain $H_{\sigma}$ accounts for every sequence that ends in a particular immune event. See Appendix A for explicit calculations; specifically (A3). As an example, consider the following,
@@
$\text{Prob}(X_{T}=\mathcal{I}^{\prime})$ is partitioned by the day of infection, which can be any day $t$ from 0 to $T-1$; this movement is indicated by the transition probability $s_{\mathcal{I}N}(t)$. Here, one is naïve before the day of infection, moves with probability 1 to the previously infected state $\mathcal{I}^{\prime}$ on the next day ($t+1$), then stays previously infected thereafter up to and including day $T$. Note that $\text{Prob}(X_{T}=N)=q_{N}(T)$; that is, the naïve state probability equals the corresponding prevalence. Similarly, $\text{Prob}(X_{T}=\mathcal{I})=f_{\mathcal{I}}(T)$ and $\text{Prob}(X_{T}=\mathcal{V})=f_{\mathcal{V}}(T)$, so the single new infection and single new vaccination state probabilities are equal to the corresponding incidences.